### v2

# Conformal Prediction for Manifold-based Source Localization with Gaussian Processes

Abstract
We tackle the challenge of uncertainty quantification in the localization of a sound source within adverse acoustic environments. Estimating the position of the source is influenced by various factors such as noise and reverberation, leading to significant uncertainty. Quantifying this uncertainty is essential, particularly when localization outcomes impact critical decision-making processes, such as in robot audition, where the accuracy of location estimates directly influences subsequent actions. Despite this, many localization methods typically offer point estimates without quantifying the estimation uncertainty. To address this, we employ conformal prediction (CP)—a framework that delivers statistically valid prediction intervals with finite-sample guarantees, independent of the data distribution. However, commonly used Inductive CP (ICP) methods require a substantial amount of labeled data, which can be difficult to obtain in the localization setting. To mitigate this limitation, we incorporate a manifold-based localization method using Gaussian process regression (GPR), with an efficient Transductive CP (TCP) technique specifically designed for GPR. We demonstrate that our method generates statistically valid uncertainty intervals across different acoustic conditions.
Index Terms:
Sound source localization (SSL), conformal prediction (CP), GPR, manifold learning.
I Introduction
Sound source localization (SSL) is essential for various audio applications, including automated camera steering [huang2000passive], speech enhancement and separation [mandel2010model], and robot audition [rascon2017localization]. Numerous classical SSL methods have been developed over the last decades, including Multiple Signal Classification (MUSIC) [schmidt1986multiple], Estimation of Signal Parameters via Rotational Invariance Techniques (ESPRIT) [roy1989esprit], Time-Difference-of-Arrival (TDOA)-based approaches [brandstein1997closed, schau1987passive], Steered Response Power (SRP) techniques [dibiase2000high] and clustering of time-difference-of-arrival (TDOA) estimates across time frequency bins [mandel2010model, schwartz2013speaker]. More recently, a variety of deep learning techniques have also been introduced [opochinsky2021deep, grinstein2023neural, grumiaux2022survey]. However, errors are unavoidable in any localization method and can significantly affect decision-making or other systems that rely on these outcomes. For example, robots and drones depend on precise location estimates to plan their next actions, while an automated camera may need to zoom out when location estimates are uncertain. Therefore, it is essential to accurately quantify the uncertainty in localization estimates.
Conformal prediction (CP) is a versatile framework for generating prediction intervals with guaranteed coverage, ensuring that the true value falls within the PIs with a user-specified probability [vovk2005algorithmic]. Recently, CP has been applied to source localization in a few studies. Previous work has explored its implementation with various uncertainty measures, such as Monte Carlo dropout, model ensembles, and quantile regression [khurjekar2023uncertainty]. In the context of multi-source direction-of-arrival (DOA) estimation, a Gaussian mixture model was employed to parameterize the conditional multi-source DOA distribution, and CP was used to derive statistical PI from the mixture model outputs [khurjekar2024multi]. However, these studies primarily focused on synthetic anechoic data and did not incorporate real speech signals. Additionally, they employ an Inductive CP (ICP) approach, which relies heavily on a substantial amount of held-out calibration data in addition to the training data.
One significant challenge in source localization is the difficulty of obtaining large and diverse labeled data that encompass various acoustic conditions, room settings, and array configurations. Most existing deep learning methods rely on simulated data from simplified room models, which may not accurately represent real-world conditions. Additionally, CP methods rely on the crucial assumption of data exchangeability, meaning that the joint distribution of calibration and test data should remain consistent under any permutation. However, this assumption is violated when calibration and test data come from different settings, making calibration on simulated data potentially unreliable when applied to real-world test data.
To address these challenges, we adopt a semi-supervised approach for source localization [laufer2016multiple, laufer2020data]. Our approach assumes access to a specific room where we can collect unlabeled data (i.e., measurements from unknown source locations) along with a small set of labeled samples with known source positions. To estimate the source positions, we employ Gaussian process regression (GPR) using a specialized manifold-based kernel that captures the geometric structure of the acoustic features, leveraging both labeled and unlabeled data. We then integrate this method with a TCP framework, which does not require additional held-out calibration data and can be efficiently applied to GPR [papadopoulos2024guaranteed]. We extensively evaluate our approach under different simulated acoustic conditions, demonstrating its effectiveness in producing valid PIs.
II Guaranteed coverage prediction intervals for Source Localization
II-A Problem Formulation
We consider a single source within a confined region surrounded by $M$ microphone nodes, each comprising two microphones. The source, positioned at $\mathbf{p}=[p_{x},p_{y}]^{T}$, emits a signal $s(k,l)$, represented in the short-time Fourier transform (STFT) domain. Here, $k\in\{1,\ldots,K\}$ denotes the frequency index and $l\in\{1,\ldots,L\}$ denotes the time index. The signal measured by the $i$-th microphone of the $m$-th node, is denoted by $x_{i,m}(k,l)$, and can be expressed as:
$x_{i,m}(k,l)=a_{i,m}(k;\mathbf{p})s(k,l)+u_{i,m}(k,l),$ (1)
where $a_{i,m}(k;\mathbf{p})$ represents the acoustic transfer function between the source at position $\mathbf{p}$ and the $i$-th microphone of the $m$-th array, and $u_{i,m}(k,l)$ denotes an additive sensor noise.
We extract an acoustic feature vector based on the relative transfer function (RTF), defined as $h(k;\mathbf{p})=\frac{a_{2,m}(k;\mathbf{p})}{a_{1,m}(k;\mathbf{p})}$, which can be estimated by:
$\bar{h}_{m}(k)=\frac{\hat{S}_{m,21}(k)}{\hat{S}_{m,11}(k)},$ (2)
where $\hat{S}_{m,21}(k)$ and $\hat{S}_{m,11}(k)$, represents the estimated cross-spectrum between microphones and the spectrum of the first microphone, respectively, when considering the measurements of the $m$-th node. We focus on a pre-defined frequency range $k_{1},\ldots,k_{F}$, where reliable RTF estimates can be expected, and define the RTF vector $\mathbf{h}^{m}=[\bar{h}_{m}(k_{1}),\ldots,\bar{h}_{m}(k_{F})]^{T}$ associated with the $m$th node. We assume that the RTF samples of the $m$-th node reside on a low dimensional manifold, denoted as $\mathcal{M}_{m}$. This assumption is based on the fact that in a given environment, where both the acoustic conditions and microphone positions remain relatively constant, the primary variable affecting the RTF samples is the change in source position [laufer2015study, laufer2020data]. Since each node has a different perspective, it will exhibit distinct relationships between the RTF samples, resulting in a unique manifold structure for each node. Finally, we define an aggregated RTF vector for all nodes by: $\mathbf{h}=[\mathbf{h}^{1},\ldots,\mathbf{h}^{m}]^{T}$.
We assume access to a training set which consists of $n_{L}$ labeled samples associated with corresponding source positions $\{\mathbf{h}_{i},\mathbf{p}_{i}\}_{i=1}^{n_{L}}$, and $n_{U}$ unlabeled samples from unknown locations $\{\mathbf{h}_{j}\}_{j=n_{L}+1}^{n}$, where $n=n_{L}+n_{U}$. Given a new test sample $\mathbf{h}_{t}$, we would like to localize the source and produce a prediction interval $\Gamma^{\delta}(\mathbf{h}_{t})$ such that:
$\mathbb{P}\left(p_{t,c}\in\Gamma^{\delta}(\mathbf{h}_{t})\right)\geq 1-\delta,$ (3)
where $c\in\{x,y\}$ and $\delta$ is a user-defined confidence level.
II-B Multiple Manifold Gaussian Process
To localize a source, we first define a mapping function associated with the $m$-th node, $f_{m}:\mathcal{M}_{m}\rightarrow\mathbb{R}$, which maps the RTF sample $\mathbf{h}^{m}\in\mathcal{M}_{m}$ to the corresponding $x$ or $y$ coordinate of the source position. Let $f^{m}_{i}\equiv f^{m}(\mathbf{h}^{m}_{i})$ denote the position evaluated by the function $f^{m}$ for the RTF sample $\mathbf{h}^{m}_{i}$, where $i$ is a sample index. Note that the mapping is applied independently to each coordinate, with the coordinate index omitted for brevity. We assume that $f^{m}$ follows a zero-mean Gaussian process (GP) with a manifold-based covariance function that quantifies the relation between samples by comparing their proximity to other samples on the manifold:
$\textrm{cov}(\mathbf{h}^{m}_{i},\mathbf{h}^{m}_{j})\equiv\tilde{k}(\mathbf{h}^{m}_{i},\mathbf{h}^{m}_{j})=\sum_{r=1}^{n}k^{m}(\mathbf{h}^{m}_{i},\mathbf{h}^{m}_{r})k^{m}(\mathbf{h}^{m}_{j},\mathbf{h}^{m}_{r}),$ (4)
where $k_{m}:\mathcal{M}_{m}\times\mathcal{M}_{m}\rightarrow\mathbb{R}$ is a standard kernel function. We use a Gaussian kernel $k_{m}(\mathbf{h}^{m}_{i},\mathbf{h}^{m}_{j})=\exp\left(-\frac{\|\mathbf{h}^{m}_{i}-\mathbf{h}^{m}_{j}\|^{2}}{\sigma_{m}}\right)$, with a scaling factor $\sigma_{m}$. In addition, we assume that the per-node processes are jointly Gaussian, with covariance:
$\textrm{cov}(\mathbf{h}^{m}_{i},\mathbf{h}^{q}_{j})\equiv\tilde{k}(\mathbf{h}^{m}_{i},\mathbf{h}^{q}_{r})=\sum_{r=1}^{n}k^{m}(\mathbf{h}^{m}_{i},\mathbf{h}^{m}_{r})k^{q}(\mathbf{h}^{q}_{j},\mathbf{h}^{q}_{r}).$ (5)
To fuse the information from the different nodes, we define a unified mapping function $f:\left(\bigcup_{m=1}^{M}\mathcal{M}_{m}\right)\rightarrow\mathbb{R}$, which associates an aggregated RTF sample $\mathbf{h}_{i}$ with a specific coordinate of the source position $f_{i}\equiv f(\mathbf{h}_{i})$. We define $f$ by the mean of the GPs of all nodes, forming a multiple-manifold GP (MMGP). Thus, each position $f_{i}$ can be written as:
$f_{i}=\frac{1}{M}\left(f^{1}_{i}+f^{2}_{i}+\cdots+f^{M}_{i}\right).$ (6)
Since the node-wise processes are jointly Gaussian, the combined process $f$ is also Gaussian with zero mean and covariance given by:
$\text{cov}(f_{i},f_{j})$ $\equiv\tilde{k}(\mathbf{h}_{i},\mathbf{h}_{j})=\frac{1}{M^{2}}\text{cov}\left(\sum_{m=1}^{M}f^{m}_{i},\sum_{g=1}^{M}f^{g}_{j}\right)$
$=\frac{1}{M^{2}}\sum_{m=1}^{M}\sum_{g=1}^{M}\text{cov}(f^{m}_{i},f^{g}_{j})$
$=\frac{1}{M^{2}}\sum_{r=1}^{n}\sum_{m,g=1}^{M}k_{m}(\mathbf{h}^{m}_{i},\mathbf{h}^{m}_{r})k_{g}(\mathbf{h}^{g}_{j},\mathbf{h}^{g}_{r}).$ (7)
This covariance function introduces a manifold-based kernel that incorporates the relations between all training points and the perspectives from different nodes.
We assume that the measured source positions satisfy a noisy observation model given by:
$p_{i}=f(\mathbf{h}_{i})+\mathbf{\epsilon}_{i},\quad i=1,\ldots,n_{L},$ (8)
where $\mathbf{\epsilon}_{i}\sim\mathcal{N}(0,\sigma_{p}^{2})$ are i.i.d. Gaussian noise terms, independent of $f(\mathbf{h}_{i})$. These noise terms accounts for uncertainties due to imperfect measurements of the source positions during the acquisition of the labeled set. To localize a new test sample $\mathbf{h}_{t}$ from an unknown source position, we use an estimator based on the posterior probability $\mathbb{P}(f(\mathbf{h}_{t})|\{\mathbf{h}_{i},p_{i}\}_{i=1}^{n_{L}},\{\mathbf{h}_{i}\}_{i=n_{L}+1}^{n})$, which is also Gaussian and is defined by the following mean and variance:
$\hat{p}_{t}=\tilde{\mathbf{K}}_{L,t}^{T}\left(\tilde{\mathbf{K}}_{L}+\sigma_{p}^{2}\mathbf{I}_{n_{L}}\right)^{-1}\mathbf{p}_{L},$ (9)
$\text{Var}(\hat{p}_{t})=\tilde{\mathbf{K}}_{t}-\tilde{\mathbf{K}}_{L,t}^{T}\left(\tilde{\mathbf{K}}_{L}+\sigma_{p}^{2}\mathbf{I}_{n_{L}}\right)^{-1}\tilde{\mathbf{K}}_{L,t},$ (10)
where $\tilde{\mathbf{K}}_{L}$ is an $n_{L}\times n_{L}$ covariance matrix defined over the labeled samples, $\tilde{\mathbf{K}}_{L,t}$ is an $n_{L}\times 1$ covariance vector between the labelled samples and $f(\mathbf{h}_{t})$, $\tilde{\mathbf{K}}_{t}$ is the variance of $f(\mathbf{h}_{t})$, and $\mathbf{I}_{n_{L}}$ is the $n_{L}\times n_{L}$ identity matrix. Note that although the variance (10) reflects estimation uncertainty, if the model is not well-specified (e.g. wrong GP model hyperparameters or likelihood [assael2014heteroscedastic, griffiths2021achieving]), these uncertainty estimates can become misleading. To mitigate this, we employ CP to produce valid PIs, even when the model is completely misspecified.
II-C Conformal Prediction with Gaussian Process Regression
CP [vovk2005algorithmic] is a framework that constructs prediction intervals (PIs) with a guaranteed coverage probability, defined by the user. There are two primary approaches to CP: Inductive CP (ICP)[papadopoulos2002inductive] and Transductive CP (TCP)[Vovk2015TransductiveCP]. ICP enhances computational efficiency by dividing the dataset into separate training and calibration subsets, while TCP leverages the entire dataset for both tasks, resulting in higher statistical efficiency. However, in TCP, the model must be re-computed for each test sample and every candidate label, which can be computationally expensive. Consequently, TCP is typically applied to specific models [nouretdinov2001ridge, papadopoulos2011regression], where this re-computation remains feasible. Recently, GPR-CP [papadopoulos2024guaranteed],was introduced, combining GPR with TCP, to generate PIs with guaranteed confidence levels while maintaining statistical efficiency. We will integrate this approach into the manifold-based GPR localization method.
The first step in any CP procedure is to define an appropriate conformity score, which measures how well a new data sample and its candidate label, $(\mathbf{h}_{t},\tilde{p}_{t})$, conform to the patterns in the existing dataset. A leave-one-out (LOO) nonconformity score is defined as:
$\alpha_{i}(\tilde{p}_{t})=\lvert p_{i}-\hat{p}_{-i}(\tilde{p}_{t})\rvert,\quad i=1,\ldots n_{L},$ (11)
where $\hat{p}_{-i}(\tilde{p}_{t})$ represents the estimator for the $i$-th sample given all other labelled samples, as well as the test sample, associated with a candidate label $\tilde{p}_{t}$. We will define $\hat{p}_{-i}$ more formally later in the context of GPR. Similarly, the nonconformity score of the test sample is $\alpha_{t}(\tilde{p}_{t})=\lvert\tilde{p}_{t}-\hat{p}_{t}\rvert$, where $\hat{p}_{t}$ is computed using (9) based on all labeled samples.
To statistically quantify how unusual the candidate label $\tilde{p}_{t}$ is, CP computes a p-value for $\tilde{p}_{t}$ by comparing its nonconformity score to those of the existing data samples:
$\mathbb{P}_{\textrm{val}}\left(\tilde{p}_{t}\right)=\frac{\#\{i=1,\dots,n_{L}:\alpha_{i}(\tilde{p}_{t})\geq\alpha_{t}(\tilde{p}_{t})\}+1}{n_{L}+1}.$ (12)
Accordingly, we construct a PI $\Gamma_{t}^{\delta}(\mathbf{h}_{t})$ as:
$\Gamma_{t}^{\delta}(\mathbf{h}_{t})=\left\{\tilde{p}_{t}:\mathbb{P}_{\textrm{val}}\left(\tilde{p}_{t}\right)>\delta\right\},$ (13)
which has a valid finite-sample coverage, as stated in the following theorem.
Theorem 1
(CP coverage guarantee [vovk2005algorithmic]). Suppose $(\mathbf{h}_{1},p_{1}),\ldots,(\mathbf{h}_{n_{L}},p_{n_{L}}),(\mathbf{h}_{t},p_{t})$ are exchangeable, then the PI (13) satisfies:
$\mathbb{P}\left(p_{t}\in\Gamma_{t}^{\delta}(\mathbf{h}_{t})\right)\geq 1-\delta.$ (14)
In principal, the computation of (11) requires re-estimating the model for each held-out combination and every candidate test position. Below, we outline how the PIs can be efficiently computed with a GPR as the underlying model [papadopoulos2024guaranteed]. We utilize the fact that the LOO estimator for GPR can be written as [rasmussen2006gaussian]:
$\hat{p}_{-i}=p_{i}-\frac{[(\mathbf{K}^{*}+\sigma_{p}^{2}\mathbf{I}_{n_{L}})^{-1}\mathbf{p}^{*}]_{i}}{[(\mathbf{K}^{*}+\sigma_{p}^{2}\mathbf{I}_{n_{L}})^{-1}]_{ii}},$ (15)
$\text{Var}(\hat{p}_{-i})=\frac{1}{[(\mathbf{K}^{*}+\sigma_{p}^{2}\mathbf{I}_{n_{L}})^{-1}]_{ii}},$ (16)
with $\mathbf{K}^{*}\in\mathbb{R}^{(n_{L}+1)\times(n_{L}+1)}$ formed by the kernel $\tilde{k}(\mathbf{h}_{i},\mathbf{h}_{j})$ in (II-B) evaluated over $\mathbf{h}_{j},\mathbf{h}_{j}\in\{\mathbf{h}_{i}\}_{i=1}^{n_{L}}\cup\mathbf{h}_{t}$, and $\mathbf{p}^{*}=[p_{1},\ldots,p_{n_{L}},\tilde{p}_{t}]^{T}$. Defining $\hat{\mathbf{p}}_{\textrm{LOO}}=[\hat{p}_{-1},\ldots,\hat{p}_{-n_{L}},\hat{p}_{t}]^{T}$, the vector of nonconformity scores $\Delta\equiv[\alpha_{1}(\tilde{p}_{t}),\ldots,\alpha_{n_{L}}(\tilde{p}_{t}),\alpha_{t}(\tilde{p}_{t})]^{T}=\lvert\mathbf{p}^{*}-\hat{\mathbf{p}}_{\textrm{LOO}}\rvert$ can be computed as:
$\lvert\Delta\rvert=\big{\lvert}(\mathbf{K}^{*}+\sigma_{p}^{2}\mathbf{I}_{n_{L}+1})^{-1}\mathbf{p}^{*}\,./\text{diag}((\mathbf{K}^{*}+\sigma_{p}^{2}\mathbf{I}_{n_{L}+1})^{-1})\big{\rvert},$ (17)
where ././. / denotes element-wise division. We decompose the vector $\mathbf{p}^{*}$ as $\mathbf{p}^{*}=\mathbf{p}_{a}+\mathbf{p}_{b}=(p_{1},\ldots,p_{n_{L}},0)^{T}+(0,\ldots,0,\hat{p}_{t})^{T}$. Thus, (17) takes the form $\lvert\mathbf{a}+\tilde{p}_{t}\cdot\mathbf{b}\rvert$ where:
$\mathbf{a}=(\mathbf{K}^{*}+\sigma_{p}^{2}\mathbf{I}_{n_{L}+1})^{-1}\mathbf{p}_{a}\,./\text{diag}((\mathbf{K}^{*}+\sigma_{p}^{2}\mathbf{I}_{n_{L}+1})^{-1}),$ (18)
$\mathbf{b}=(\mathbf{K}^{*}+\sigma_{p}^{2}\mathbf{I}_{n_{L}+1})^{-1}\mathbf{p}_{b}\,./\text{diag}((\mathbf{K}^{*}+\sigma_{p}^{2}\mathbf{I}_{n_{L}+1})^{-1}).$ (19)
Hence, the nonconformity scores $\lvert\Delta\rvert=\lvert\mathbf{a}+\tilde{p}_{t}\cdot\mathbf{b}\rvert$ are piecewise linear with respect to $\tilde{p}_{t}$ for each $\Delta_{i}$. Furthermore, since $\mathbb{P}_{\textrm{val}}(\tilde{p}_{t})$ changes only when $\alpha_{i}(\tilde{p}_{t})=\alpha_{t}(\tilde{p}_{t})$, it eliminates the need to evaluate infinitely many candidates, thereby resulting in a feasible prediction algorithm. Specifically, for computing $\mathbb{P}_{\textrm{val}}(\tilde{p}_{t})$, we construct finite-sets for every labelled sample:
$S_{i}$ =$=\left\{\tilde{p}_{t}:\alpha_{i}(\tilde{p}_{t})\geq\alpha_{t}(\tilde{p}_{t})\right\}\>\>\>\>\>\>\>\>,i=1,\dots,n_{L}$
=$=\left\{\tilde{p}_{t}:|a_{i}+\tilde{p}_{t}\cdot b_{i}|\geq|a_{n_{L}+1}+\tilde{p}_{t}\cdot b_{n_{L}+1}|\right\}.$ (20)
Then, the p-value (12) takes the form:
$\mathbb{P}_{\textrm{val}}(\tilde{p}_{t})=\frac{\#\{i=1,\dots,n_{L}:\tilde{p}_{t}\in S_{i}\}+1}{n_{L}+1},$ (21)
and the interval $\Gamma_{t}^{\delta}(\mathbf{h}_{t})$ contains all candidates for which $\mathbb{P}_{\textrm{val}}(\tilde{p}_{t})>\delta$, as defined in (13). An efficient algorithm for finding all relevant intervals appears in [papadopoulos2024guaranteed, nouretdinov2001ridge].
In addition, the nonconformity scores can be normalized to obtain more precise PIs, which are tighter for inputs that are easier to predict and wider for inputs that are more difficult to predict [papadopoulos2024guaranteed]. We use the LOO-variance (16) for normalization:
$\alpha_{i}(\tilde{p}_{t})=\left\lvert\frac{p_{i}-\hat{p}_{-i}}{\sqrt[\gamma]{\text{Var}(\hat{p}_{-i})}}\right\rvert=\left\lvert\frac{p_{i}-\hat{p}_{-i}}{[(\mathbf{K}^{*}+\sigma_{p}^{2}\mathbf{I}_{n_{L}+1})^{-1}]_{ii}^{-\frac{1}{\gamma}}}\right\rvert,$ (22)
where $\gamma$ controls the sensitivity to changes in the variance. Accordingly, the nonconformity scores (17) take the form:
$\lvert\Delta\rvert=\big{\lvert}(\mathbf{K}^{*}+\sigma_{p}^{2}\mathbf{I}_{n_{L}+1})^{-1}\mathbf{p}^{*}\,./\text{diag}((\mathbf{K}^{*}+\sigma_{p}^{2}\mathbf{I}_{n_{L}+1})^{-1})^{1-\frac{1}{\gamma}}\big{\rvert}.$
III Experiments and Results
III-A Experimental settings
Figure 1: The simulated room setup.
We simulated $5.2\times 6.2\times 3.5$m rooms with reverberation times of $300$ ms and $700$ ms. Five ($M=5$) microphone pairs were placed around the room’s walls. The source positions were confined to a $2\times 2$m squared region (see Fig. 1). We generated $n_{L}=225$ labeled samples, equally spaced with $0.133$m resolution, $n_{U}=100$ unlabeled samples and $n_{T}=200$ test samples, both randomly sampled within the same region. Simulated room impulse responses [habets2006rirgenerator] were convolved with white Gaussian noise signals to produce the labeled measurements, while speech utterances from the LibriSpeech dev-set [panayotov2015librispeech] were used for the unlabeled and test samples. White Gaussian noise was added to all microphone signals, with an SNR of $15$dB. The signals were analyzed using an STFT with $1024$ frequency bins and $75\%$ frame overlap. The frequency range $150-1500$Hz was used in construction of the RTF vectors. GPyTorch [gardner2018gpytorch] was employed for optimizing the model’s parameters $\Theta=\{\sigma_{1},\ldots,\sigma_{M},\sigma_{p}\}$. We set $\gamma=8$ for the score normalization (22).
III-B Evaluation
We compare the performance of the proposed approach (GPR-CP) with the standard GPR, where PIs are derived directly from the estimated variance (10). Note that in case of model misspecification, the PIs produced by the GPR do not have coverage guarantees. We examine two coverage levels: $95$% ($\delta=0.05$) and $99$% ($\delta=0.95$). We assess whether the empirical coverage rate of the PIs matches the nominal coverage rate. Moreover we evaluate the tightness of the PI by their mean width, where narrower PIs are desirable, provided they satisfy the required coverage levels. We repeat each experiment 10 times, using different random unlabeled and test positions, and report the average results across these trials.
III-C Results
The empirical coverage and mean PI width for $300$ ms and $700$ ms RTs are presented in Figs. 2 and  3, respectively. Due to space constraints, we present the results along the x-axis, though similar trends were observed for the y-axis. Noticeably, GPR-CP obtains the nominal coverage in all cases, while GPR losses coverage at moderate reverberation. In addition, for high-reverberation, where the coverage of both methods is comparable, GP-CP produces smaller intervals. For $T_{60}=300$ ms and $\delta=0.05$, both GPR-CP and GPR exhibit almost similar interval widths, even though GPR does not satisfy the coverage. For $T_{60}=300$ ms and $\delta=0.01$, we observe wider PIs for GPR-CP compared to GPR, which is acceptable, since the coverage requirement is fulfilled. Fig. 4 depicts the PIs produced for the X and Y coordinates as a function of the true position, with red x-marks denoting the labeled positions. Here too, we observe that GPR-CP often produces smaller intervals compared to those of GPR. Evidently, the width of the PIs on each coordinate is dictated by how well-distributed the microphones are. Since there is a single node placed across the Y-axis (compared to $4$ nodes for the X-axis), it makes predictions less certain, leading to wider intervals. In addition, it can be seen that the interval width is smaller near labeled samples, as expected.
Figure 2: Coverage (left) and mean width (right) - $T_{60}=300$ ms.
Figure 3: Coverage (left) and mean width (right) - $T_{60}=700$ ms.
Figure 4: PI width at $15$dB for $95\%$ coverage at $T_{60}=700$ ms. Left panel: X-coordinate. Right panel: Y-coordinate.
IV Conclusions
We propose a method for uncertainty quantification in estimating speaker location in noisy and reverberant environments. Our approach utilizes a small set of labeled data, complemented by additional unlabeled data from unknown source locations. The core of our method is a GPR model with a manifold-based kernel function, integrated with a CP framework, tailored specifically for GPR. This combination enables the construction of valid prediction intervals with guaranteed coverage, which can be computed efficiently. Results are demonstrated across different simulated acoustic conditions.